# Magnet: Augmenting Generative Decoders with Representation Learning and Infilling Capabilities

### v1

@@ Abstract
While originally designed for unidirectional generative modeling, decoder-only large language models (LLMs) are increasingly being adapted for bidirectional modeling. However, unidirectional and bidirectional models are typically trained separately with distinct objectives (generation and representation learning, respectively). This separation overlooks the opportunity for developing a more versatile language model and for these objectives to complement each other. In this work, we introduce Magnet, an adaptation of decoder-only LLMs that enhances their ability to generate robust representations and infill missing text spans, while preserving their knowledge and text generation capabilities. Magnet employs three self-supervised training objectives and introduces an attention mechanism that combines bidirectional and causal attention, enabling unified training across all objectives. Our results demonstrate that LLMs adapted with Magnet (1) surpass strong text encoders on token-level and sentence-level representation learning tasks, (2) generate contextually appropriate text infills by leveraging future context, (3) retain the ability for open-ended text generation without exhibiting repetition problem, and (4) preserve the knowledge gained by the LLM during pretraining.
@@ 1 Introduction
Language models are computational models designed to understand and generate human language. They have transformed natural language processing, powering applications such as text annotation, machine translation, summarization, speech recognition, and dialogue systems. Traditionally, language models are categorized into three main types: (1) Encoder-only models (Devlin et al., 2019; Liu et al., 2019; Lan et al., 2019; Sanh et al., 2019), which focus on encoding the input into fixed-dimensional representations and excel in tasks like sentiment analysis (sentence-level classification) and named entity recognition (token-level classification). (2) Decoder-only models (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023a; b; Scao et al., 2022; Penedo et al., 2023), which are adept at generating coherent text, thereby specializing in tasks like creative content generation and dialogue systems. (3) Encoder-decoder models (Raffel et al., 2019; Lewis et al., 2019), wherein an encoder understands the input and a decoder generates the corresponding output, making this architecture suitable for tasks like machine translation and summarization.
Recently, the NLP community has increasingly embraced decoder-only architecture (Large Language Models or LLMs) due to their efficient training and scalability to larger datasets, resulting in enhanced performance across various tasks. However, since these models are trained using causal attention and lack bidirectional context, they are less suitable for tasks like (1) sentiment analysis and named entity recognition, which require understanding contextual representations of sentences or words, and (2) text infilling, where predicting missing text spans must maintain coherence with subsequent context. With the goal of leveraging the scalable decoder-only LLMs, some recent efforts (BehnamGhader et al., 2024; Li & Li, 2023; Li et al., 2023; Duki’c & vSnajder, 2024; Du et al., 2021; Donahue et al., 2020) have aimed to adapt decoder-only LLMs for these tasks. However, as illustrated in Figure 1, methods that enhance LLMs for text infilling do not enable them to function as effective text encoders, while approaches focused on representation learning hinder their generative capabilities.
In this work, we propose Magnet (Modified Attention for Generation and Encoding of Text), an adaptation of decoder-only LLMs for (1) generating robust sentence-level and token-level representations, (2) infilling missing spans of text while preserving coherence with bidirectional context, and (3) performing open-ended text generation without excessive repetition of words or phrases. To achieve this, we use three self-supervised training objectives: (1) a masked modeling objective to learn token-level representations, (2) a contrastive objective to learn sentence-level representations, and (3) a missing-span generation objective to infill text and retain generative capabilities. To facilitate simultaneous training across all these objectives, we deploy a specially crafted attention mask that combines bidirectional and causal attention.
Without any model-specific design or loss of generality, we apply Magnet to LLaMA-2-7B (Touvron et al., 2023b). We demonstrate that the proposed method requires simple modification and fine-tuning of an off-the-shelf LLM to augment it with representation learning and infilling capabilities. Our results show that Magnet-adapted LLaMA-2-7B outperforms other methods that adapt the same model for token-level and sentence-level representation learning tasks [^1] It is to be noted that these other methods exclusively adapt the model for learning representations, while Magnet trains for other objectives as well.. We also show that Magnet significantly improves the infilling capability of the LLM by enabling it to consider the bidirectional context. Further, we analyze the repetition problem in text generated by text encoders and demonstrate that Magnet-adapted models are significantly better at open-ended text generation than other text encoders. Lastly, we show that LLMs adapted using Magnet retains the knowledge gained during pretraining.
@@ 2 Related Works
Representation Learning.  Text representation learning focuses on understanding contextual relationships within sentences. Traditionally, encoder models dominated this field due to their bidirectional context modeling, using masked language modeling for token-level representations (Devlin et al., 2019; Liu et al., 2019; He et al., 2020; Clark et al., 2020; He et al., 2021) and special tokens with similarity-based optimization for sentence-level understanding (Gunel et al., 2020; Reimers & Gurevych, 2019; Wu et al., 2020; Carlsson et al., 2021; Gao et al., 2021; Wei et al., 2020). Recent work has explored adapting decoder-only LLMs for representation learning, either by using last-token or mean-pooled representations (Neelakantan et al., 2022; Wang et al., 2023), or by fine-tuning with masked modeling (BehnamGhader et al., 2024) or label supervision (Li et al., 2023; Duki’c & vSnajder, 2024). While some approaches modify the decoder’s causal attention to be bidirectional (BehnamGhader et al., 2024; Muennighoff et al., 2024; Li & Li, 2023; Duki’c & vSnajder, 2024), this often compromises the model’s text generation abilities. In contrast, Magnet employs a hybrid attention mechanism that combines causal and bidirectional attention, enabling both robust representation learning and preserved generation capabilities.
Text Infilling.  Text infilling requires considering both left and right context when generating text in the middle of a sequence. Encoder-decoder models (Raffel et al., 2019; Lewis et al., 2019; Kalinsky et al., 2023) can handle this task by encoding available context and decoding infilled text. Other approaches have extended masked language modeling to perform span infilling (Joshi et al., 2019; Shen et al., 2023; 2020). Decoder-only models have also been adapted for infilling through various strategies: training models to directly fill marked blanks (Donahue et al., 2020; Du et al., 2021), rearranging training examples to align with infilling objectives (Bavarian et al., 2022; Yang et al., 2019; Aghajanyan et al., 2022; Fried et al., 2022), or using dual generation from both ends of a sentence until convergence (Nguyen et al., 2023; Serdyuk et al., 2017). However, while these approaches successfully enhance LLMs with infilling capabilities, none have attempted to simultaneously equip them with both infilling and representation learning abilities, as done by Magnet.
@@ 3 Method
Decoder-only models, based on the Transformer architecture (Vaswani et al., 2017), process input sequences through successive blocks of multi-head self-attention, feed-forward networks, and layer normalization. The self-attention mechanism converts the input $\mathbf{x}{}\in\mathbb{R}^{l\times d}$ into queries $\mathbf{Q}{}$, keys $\mathbf{K}{}$, and values $\mathbf{V}{}$ using a linear projections, and computes attention using the formula:
$\texttt{Attn}_{i}(\mathbf{Q}{},\mathbf{K}{},\mathbf{V}{})=\texttt{softmax}\left(\frac{\mathbf{Q}{}\mathbf{K}{}^{\text{T}}+\mathbf{M}{}}{\sqrt{d_{k}}}\right)\mathbf{V}{}$ (1)
where $\texttt{Attn}_{i}$ is the $i^{th}$ head of the multi-head self-attention, $d_{k}$ represents the dimensionality of the keys/queries, and $\mathbf{M}{}$ represents the causal mask.
Magnet seeks to update the causal attention mechanism of an LLM by incorporating elements of bidirectionality and thereafter fine-tunes the model using unsupervised objectives. We look at the modifications to the attention mechanism in Section 3.1 and the training objectives in Section 3.2.
3.1 Modifying Attention
Magnet updates the causal attention mechanism of an LLM to introduce bidirectional capabilities within segments of the input sequence, as illustrated in Figure 2. After the input text is tokenized for the language model, we categorize each token as either context tokens or span tokens:
Context tokens.  Each context token (shown in blue in Figure 2) attends to all other context tokens within the sequence. In our implementation, the attention mask is designed with 0s at output positions corresponding to context tokens, allowing each context token to access information from every other context token. This transformation shifts the original unidirectional LLM into a bidirectional model.
Span tokens.  The span tokens (shown in green in Figure 2) represent a contiguous span of input tokens that attend to all context tokens and have causal attention among themselves. By enabling span tokens to access surrounding context, we effectively convert the original LLM into an infilling language model. Additionally, the causal attention among span tokens preserves the LLM’s generative capabilities, which could be compromised if bidirectionality is fully unlocked (see Section 4.4 for further details).
During training, an input sequence includes one or more spans of span tokens surrounded by context tokens. During inference, the attention mechanism can operate in three modes: (1) fully causal/unidirectional for open-ended text generation tasks, (2) fully bidirectional for tasks requiring representation learning, or (3) a combination of causal and bidirectional for text infilling.
@@ 3.1 Modifying Attention
Figure 2: Matrices illustrating attention masks for different types of attention mechanisms. The rows of the matrices correspond to the query tokens and the columns correspond to the key tokens. Light gray cells indicate 0, dark gray cells represent $\infty$- ∞, green marks span token positions, and blue marks context token positions. Each context token attends to every other context token, and each span token attends to all context tokens and the preceding span tokens in the same span.
@@ 3.2 Training Objectives
Magnet fine-tunes an off-the-shelf LLM using three self-supervised objectives aimed at enhancing the model’s ability to learn contextually-rich representations and autoregressively fill in missing spans of text. These objectives are illustrated in Figure 3 and discussed below.
@@ 3.2.1 Masked Next Token Prediction (MNTP)
[…]
$\mathcal{L}_{\text{MNTP}}=-\frac{1}{NL}\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{v=1}^{V}\mathbbm{1}_{\text{mask}}(l+1)\cdot(y^{(n)}_{lv}\log(\hat{y}^{(n)}_{lv}))$ (2)
where, $N$ denotes batch size, $L$ denotes the sequence length, $V$ denotes vocabulary size, $\mathbbm{1}_{\text{mask}}(l+1)$ is 1 if position $l+1$ is masked and 0 otherwise, and $y_{lv}$ and $\hat{y}_{lv}$ represent the true and predicted probabilities for the $v^{th}$ token in the vocabulary at position $l$ in the sequence. It is to be noted that this task is conducted exclusively with the context tokens.
@@ 3.2.2 Self-Supervised Contrastive Learning (SSCL)
Since LLMs are not explicitly trained to capture the entire input context and generate sentence-level representations, we employ SSCL to transform them into text encoders. The training objective is defined as follows: Given an input sequence $\mathbf{x}{}$, generate its augmented view $\mathbf{x}{}^{\mathbf{+}}$ and align their encoded representations $\mathbf{e}=f(\mathbf{x}{})$ and $\mathbf{e^{\mathbf{+}}}=f(\mathbf{x}{}^{\mathbf{+}})$ in the embedding space, while distancing them from the encodings $\mathbf{e^{\mathbf{-}}}=f(\mathbf{x}{}^{\mathbf{-}})$ of other input sequences $\mathbf{x}{}^{\mathbf{-}}$ in a training batch. Specifically, we employ paraphrasing (Damodaran, 2021) to generate augmented views of an input, and add an instruction "Given the sentence, find its representation:" to the training examples (Jiang et al., 2023). Then, we use the representations corresponding to the last token ([EOS]) of the final hidden states as the sentence encoding. Our choice of using the last token representation as the encoding is guided by the fact that Magnet optimizes simultaneously for token-level and sentence-level representations. Since the last token’s representation is not used for token-level optimization (because the representation of input token $i$ is given by output token $i-1$), this choice enables us to disentangle the two representation learning tasks during joint training. We use InfoNCE (van den Oord et al., 2018) with in-batch negatives as the loss function:
$\mathcal{L}_{\text{SSCL}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\textbf{e}_{i}\cdot\textbf{e}^{+}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\textbf{e}_{i}\cdot\textbf{e}^{-}_{j}/\tau)}$ (3)
where, $N$ represents the batch size and $\tau$ denotes the temperature for logit scaling.
@@ 3.2.3 Missing Span Generation (MSG)
MSG provides text infilling capabilities to the left-to-right autoregressive model. The task is defined as: Given a position $p$ and an input sequence $\mathbf{x}{}=(x_{1},...,x_{p},x_{q},...,x_{L})$, generate a plausible sequence of $m$ tokens $\mathbf{y}{}=(y_{1},y_{2},...,y_{m})$ that fits between $x_{p}$ and $x_{q}$. More specifically, in our training setup, this task entails predicting a span token $y_{l}$ conditioned on all context tokens in $\mathbf{x}{}$ and the preceding span tokens $x_{[1..l-1]}$. We train using categorical cross-entropy loss computed over the predicted span tokens:
$\mathcal{L}_{\text{MSG}}=-\frac{1}{N}\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{v=1}^{V}\mathbbm{1}_{\text{span}}(l)\cdot(y^{(n)}_{lv}\log(\hat{y}^{(n)}_{lv}))$ (4)
where, $N$ denotes batch size, $L$ denotes sequence length, $V$ denotes vocabulary size, $\mathbbm{1}_{\text{span}}(l)$ is 1 if the token at position $l$ is a span token and 0 otherwise, and $y_{lv}$ and $\hat{y}_{lv}$ are the true and predicted probabilities for token $v$ in the vocabulary at position $l$ in the sequence. The standard next token prediction task of LLMs can be considered as a special case of this objective, wherein all input tokens are span tokens (and the attention mechanism reduces to causal attention). Thus, a beneficial side-effect of this task is that the model retains its text generation capability while learning representations.
Figure 4: Magnet processes three views of the input using different attention mechanisms within the same LLM. The model is trained (or fine-tuned) using three self-supervised learning objectives simultaneously to augment it with the ability to generate token-level and sentence-level representations, and perform text infilling tasks, while maintaining its original left-to-right text generation capability.

### v2

@@ Abstract
While originally designed for unidirectional generative modeling, decoder-only large language models (LLMs) are increasingly being adapted for bidirectional modeling. However, unidirectional and bidirectional models are typically trained separately with distinct objectives (generation and representation learning). This separation overlooks the opportunity for developing a more versatile language model and for these objectives to complement each other. In this work, we propose Magnet, a method for adapting decoder-only LLMs to generate robust representations and infill missing text spans. Magnet employs three self-supervised training objectives and introduces an attention mechanism that combines bidirectional and causal attention, enabling unified training across all objectives. Our results demonstrate that LLMs adapted with Magnet (1) surpass strong text encoders on token-level and sentence-level representation learning tasks, (2) generate contextually appropriate text infills by leveraging past and future contexts, (3) perform open-ended text generation without excessive repetition of words or phrases, and (4) preserve the knowledge and reasoning capability gained by the LLM during pretraining.
Magnet: Augmenting Generative Decoders with Representation Learning and Infilling Capabilities
Savya Khosla12thanks: Work done during internship at Adobe Research. Correspondence to savyak2@illinois.edu., Aditi Tiwari2, Kushal Kafle1, Simon Jenni1, Handong Zhao1, John Collomosse1, Jing Shi1 1Adobe Research, 2University of Illinois Urbana-Champaign
@@ 1 Introduction
Decoder-only LLMs have gained popularity in NLP due to their efficient training and scalability. However, their reliance on causal attention restricts their effectiveness in tasks that require understanding of bidirectional context. This limitation is particularly evident in (1) representation learning tasks such as sentiment analysis and named entity recognition, where understanding the full context of sentences or words is crucial, and (2) text infilling, where filling in missing spans must ensure coherence with the surrounding text.
Some recent efforts (BehnamGhader et al., 2024; Li and Li, 2023; Li et al., 2023; Duki’c and vSnajder, 2024; Du et al., 2021; Donahue et al., 2020) have sought to adapt decoder-only LLMs for representation learning and text infilling. However, as shown in Figure 1, methods that enhance LLMs for text infilling fail to make them effective text encoders, while methods focused on representation learning diminish their generative capabilities.
In this work, we introduce Magnet (Modified Attention for Generation and Encoding of Text), a method for adapting decoder-only LLMs into more versatile language models. Magnet enables an LLM to (1) generate robust sentence-level and token-level representations, (2) infill missing text spans while maintaining coherence with bidirectional context, (3) perform open-ended text generation without excessive repetition, and (4) preserve the knowledge acquired during pretraining. In essence, Magnet equips an LLM with representation learning and infilling capabilities while preserving its generative strengths. To achieve this, we use three self-supervised training objectives: (1) a masked modeling objective to learn token-level representations, (2) a contrastive objective to learn sentence-level representations, and (3) a missing-span generation objective to infill text and retain generative capabilities. To facilitate simultaneous training across all these objectives, we deploy a specially crafted attention mask that combines bidirectional and causal attention.
Without any model-specific design, we apply Magnet to Llama-2-7B (Touvron et al., 2023). We demonstrate that the proposed method requires simple modification and fine-tuning of an off-the-shelf LLM to augment it with representation learning and infilling capabilities. Our results show that Magnet-adapted Llama-2-7B outperforms other methods that adapt the same model for token-level and sentence-level representation learning tasks [^1] It is to be noted that while these other methods adapt the model exclusively for representation learning, Magnet incorporates additional objectives, making the LLM more versatile and showcasing the advantages of unified training.. We also show that Magnet improves the infilling capability of the LLM by enabling it to consider the bidirectional context. Further, we analyze the repetition problem in text generated by models that are trained or fine-tuned to encode text and demonstrate that Magnet-adapted models are significantly better at open-ended text generation than other text encoders. Lastly, we show that Magnet preserves the knowledge and reasoning capabilities acquired by the LLM during pretraining.
@@ 2 Related Works
Representation Learning.  Text representation learning focuses on understanding contextual relationships within sentences. Traditionally, encoder models dominated this field due to their bidirectional context modeling, using masked language modeling for token-level representations (Devlin et al., 2019; Liu et al., 2019; He et al., 2020; Clark et al., 2020; He et al., 2021) and special tokens with similarity-based optimization for sentence-level understanding (Gunel et al., 2020; Reimers and Gurevych, 2019; Wu et al., 2020; Carlsson et al., 2021; Gao et al., 2021; Wei et al., 2020). Recent work has explored adapting decoder-only LLMs for text encoding through various methods, including introducing special tokens to the model’s vocabulary (Zhang et al., 2024), using last-token or mean-pooled representations (Neelakantan et al., 2022; Wang et al., 2023), or fine-tuning with masked modeling (BehnamGhader et al., 2024) or label supervision (Li et al., 2023; Duki’c and vSnajder, 2024). While some approaches modify the decoder’s causal attention to be bidirectional (BehnamGhader et al., 2024; Muennighoff et al., 2024; Li and Li, 2023; Duki’c and vSnajder, 2024; Man et al., 2024), this often compromises the model’s text generation abilities. In contrast, Magnet employs a hybrid attention mechanism that combines causal and bidirectional attention, enabling both robust representation learning and preserved generation capabilities.
Text Infilling.  Text infilling requires considering both left and right context when generating text in the middle of a sequence. Encoder-decoder models (Raffel et al., 2019; Lewis et al., 2019; Kalinsky et al., 2023) can handle this task by encoding available context and decoding infilled text. Other approaches have extended masked language modeling to perform span infilling (Joshi et al., 2019; Shen et al., 2023, 2020). Decoder-only models have also been adapted for infilling through various strategies: training models to directly fill marked blanks (Donahue et al., 2020; Du et al., 2021), rearranging training examples to align with infilling objectives (Bavarian et al., 2022; Yang et al., 2019; Aghajanyan et al., 2022; Fried et al., 2022), or using dual generation from both ends of a sentence until convergence (Nguyen et al., 2023; Serdyuk et al., 2017). However, while these approaches successfully enhance LLMs with infilling capabilities, none have attempted to simultaneously equip them with both infilling and representation learning abilities, as done by Magnet.
@@ 3 Method
Decoder-only models process input sequences through successive blocks of multi-head self-attention, feed-forward networks, and layer normalization. The self-attention mechanism converts the input $\mathbf{x}{}\in\mathbb{R}^{l\times d}$ into queries $\mathbf{Q}{}$, keys $\mathbf{K}{}$, and values $\mathbf{V}{}$ using linear projections, and computes attention using the formula:
$\texttt{Attn}_{i}(\mathbf{Q}{},\mathbf{K}{},\mathbf{V}{})=\texttt{softmax}\left(\frac{\mathbf{Q}{}\mathbf{K}{}^{\text{T}}+\mathbf{M}{}}{\sqrt{d_{k}}}\right)\mathbf{V}{}$
where $\texttt{Attn}_{i}$ is the $i^{th}$ head of the multi-head self-attention, $d_{k}$ represents the dimensionality of the keys/queries, and $\mathbf{M}{}$ represents the causal mask. This causal mask $\mathbf{M}{}$ for an autoregressive LLM is a $l\times l$ strictly upper triangular matrix, as shown in Figure 2(a), and it ensures that each token can only attend to itself and tokens that precede it.
Magnet updates the causal attention mechanism of an LLM to incorporate elements of bidirectionality and thereafter fine-tunes the model using self-supervised objectives. We look at the modifications to the attention mechanism in Section 3.1 and the training objectives in Section 3.2.
@@
Figure 2: Attention masks for different types of attention mechanisms. The rows of the matrices correspond to the query tokens and the columns correspond to the key tokens. Light gray cells indicate 0, dark gray cells represent $\infty$- ∞, green marks span token positions, and blue marks context token positions. Each context token attends to every other context token, and each span token attends to all context tokens and the preceding span tokens in the same span.
3.1 Modifying Attention
We categorize the input tokens as either context tokens or span tokens and use the attention mask shown in Figure 2.
Context tokens.  Each context token (shown in blue in Figure 2) attends to all other context tokens in the sequence. The attention mask has 0s at output positions corresponding to context tokens, allowing each context token to access information at every other context token. This transformation shifts the original unidirectional LLM into a bidirectional model.
Span tokens.  The span tokens (shown in green in Figure 2) are a contiguous span of input tokens that attend to all context tokens and have causal attention among themselves. By enabling span tokens to access surrounding context, we effectively convert the original LLM into an infilling language model. Additionally, the causal attention among span tokens preserves the LLM’s generative capabilities, which could be compromised if bidirectionality is fully unlocked (see Section 4.4 for details).
During training, an input sequence includes one or more spans of span tokens surrounded by context tokens. During inference, the attention mechanism can operate in three modes: (1) fully causal/unidirectional for open-ended text generation tasks, (2) fully bidirectional representation learning tasks, or (3) a combination of causal and bidirectional for text infilling.
Figure 4: Magnet processes three views of the input using different attention mechanisms within the same LLM. The model is trained (or fine-tuned) using three self-supervised learning objectives simultaneously to augment it with the ability to generate token-level and sentence-level representations and perform text infilling tasks, while maintaining its original left-to-right text generation capability.
@@ 3.2 Training Objectives
Magnet fine-tunes an off-the-shelf LLM using three self-supervised objectives. These objectives are illustrated in Figure 3 and discussed below.
@@ 3.2.1 Masked Next Token Prediction (MNTP)
MNTP enables the model to realize its newly enabled bidirectional attention capability. The task is defined as follows: Given an input sequence $\mathbf{x}{}=(x_{1},x_{2},...,x_{L})$, we select a fraction of the input tokens for masking and train the model to predict these masked tokens. In our setup, we find that selecting 20% of the input tokens for masking works well. Further, following Devlin et al. (2019), we replace 80% of the selected tokens with a [MASK] token, 10% with a random token from the model’s vocabulary, and leave the remaining 10% unchanged. Since LLMs are trained to predict the next token in a sequence, we use the token representations from position $l$ to predict a masked token at position $l+1$ (as shown in Figure 3a). In Appendix D, we also explore the possibility of using the standard masked token prediction (MTP) objective, where the output at token $l$ predicts the masked token at position $l$ and find that MTP performs poorly for LLMs that are trained to predict autoregressively. MNTP is optimized using categorical cross-entropy loss:
$\mathcal{L}_{\text{MNTP}}=$ $\frac{-1}{NL}\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{v=1}^{V}$
$\left(\mathbbm{1}_{\text{mask}}(l+1)\cdot(y^{(n)}_{lv}\log(\hat{y}^{(n)}_{lv}))\right)$
where $N$ denotes batch size, $L$ denotes sequence length, $V$ denotes vocabulary size, $\mathbbm{1}_{\text{mask}}(l+1)$ is 1 if position $l+1$ is masked and 0 otherwise, and $y_{lv}$ and $\hat{y}_{lv}$ represent true and predicted probabilities for $v^{th}$ token in vocabulary at position $l$ in the sequence. Note that this task is conducted exclusively with the context tokens.
@@ 3.2.2 Self-Supervised Contrastive Learning (SSCL)
Since LLMs are not explicitly trained to capture the entire input context and generate sentence-level representations, we employ SSCL to transform them into text encoders. The task is defined as follows: Given an input sequence $\mathbf{x}{}$, we generate its augmented view $\mathbf{x}{}^{\mathbf{+}}$ and align their encoded representations $\mathbf{e}=f(\mathbf{x}{})$ and $\mathbf{e^{\mathbf{+}}}=f(\mathbf{x}{}^{\mathbf{+}})$ in the embedding space, while distancing them from the encodings $\mathbf{e^{\mathbf{-}}}=f(\mathbf{x}{}^{\mathbf{-}})$ of other input sequences $\mathbf{x}{}^{\mathbf{-}}$ in a training batch. Specifically, we employ paraphrasing (Damodaran, 2021) to generate augmented views of an input, and add an instruction "Given the sentence, find its representation:" to the training examples (Jiang et al., 2023). Then, we use the output corresponding to the last token ([EOS]) of the final hidden states as the sentence encoding. Our choice of using the last token representation as the encoding is guided by the fact that Magnet optimizes simultaneously for token-level and sentence-level representations. Since the last token’s representation is not used for token-level optimization (because the representation of input token $l$ is given by output token $l-1$), this choice enables us to disentangle the two representation learning tasks during joint training. We use InfoNCE (van den Oord et al., 2018) with in-batch negatives as the loss function:
$\mathcal{L}_{\text{SSCL}}=\frac{-1}{N}\sum_{i=1}^{N}\log\frac{\exp(\textbf{e}_{i}\cdot\textbf{e}^{+}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\textbf{e}_{i}\cdot\textbf{e}^{-}_{j}/\tau)}$
where $N$ represents the batch size and $\tau$ denotes the temperature for logit scaling.
@@ 3.2.3 Missing Span Generation (MSG)
MSG provides text infilling capabilities to the left-to-right autoregressive model. The task is defined as: Given a position $p$ and an input sequence $\mathbf{x}{}=(x_{1},...,x_{p},x_{q},...,x_{L})$, generate a plausible sequence of $m$ tokens $\mathbf{y}{}=(y_{1},y_{2},...,y_{m})$ that fits between $x_{p}$ and $x_{q}$. More specifically, in our training setup, this task entails predicting a span token $y_{l}$ conditioned on all context tokens in $\mathbf{x}{}$ and the preceding span tokens $y_{[1..l-1]}$. We train using categorical cross-entropy loss computed over the predicted span tokens:
$\mathcal{L}_{\text{MSG}}=\frac{-1}{N}\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{v=1}^{V}\mathbbm{1}_{\text{span}}(l)\cdot(y^{(n)}_{lv}\log(\hat{y}^{(n)}_{lv}))$
where $N$ denotes batch size, $L$ denotes sequence length, $V$ denotes vocabulary size, $\mathbbm{1}_{\text{span}}(l)$ is 1 if the token at position $l$ is a span token and 0 otherwise, and $y_{lv}$ and $\hat{y}_{lv}$ are the true and predicted probabilities for token $v$ in the vocabulary at position $l$ in the sequence. The standard next token prediction task of LLMs can be considered as a special case of this objective, wherein all input tokens are span tokens (and the attention mechanism reduces to causal attention). Thus, a beneficial side effect of this task is that the model retains its text generation capability while learning bidirectional representations.